# Rethinking the effects of data contamination in Code Intelligence

### v2

@@ II-E3 RQ3: Does Unpaired Contamination Lead to Performance Overestimation?
As for code generation and summarization tasks, it is challenging to find an independent NLD that coincidentally describes a code functionality but is not attached to a code snippet, because code and comments are normally closely attached in the wild. Therefore, we ignore the above two tasks for LLMs in the unpaired contamination setting.
II-E4 RQ4: Does Paired Contamination Lead to Performance Overestimation?
In this RQ, PLMs are still experimented with code translation and generation tasks. As for LLMs, we carry out the code generation task for exploration.
Pre-trained Language Models: To conduct code translation (JavaC#) experiments with PLMs in this setting, we still follow the procedures of RQ1 to establish the uncontaminated control group. As for the contaminated experimental group, we concatenate the input and output portions of CodeTrans’s testing set with a separator token <SEP> and inject these concatenated samples into the pre-training corpora of PLMs. We use <SEP> to separate different PLs, as it is a frequently-used method for bi-modal pre-training in practice [7, 12, 40]. Similarly, we also remove the corresponding number of samples from each pre-training corpus to eliminate the influence of pre-training amount differences. As for the code generation task (NLJava), we follow the same procedures for experiments.
Large Language Models: To construct paired contamination experiments on the code generation task (NLJava), we first use tree-sitter to identify and extract Java functions from both pre-training corpora. To curate samples containing both NLDs and code in order, we examine each sample to determine whether its NL description appears closely in front of the function signature. Afterwards, we perform manual filtering to select samples where the NL descriptions are accurate and informative enough to be used for code generation. The uncontaminated counterpart is prepared following the same perturbation procedures as RQ1.
As for code summarization, the inputs are code snippets while the outputs are NLDs. Although code-comment pairs are prevalent in practice, their comments are always in front of the corresponding code, which is opposite to the task definition of code summarization. Thus, we find difficulty curating such paired samples from LLMs’ pre-training corpora and ignore this task here. Moreover, in practice, programs of different PLs are very uncommon to be written in the same source files, even if they carry identical logic. Thus, it is hard to find two programs with equivalent logic pre-trained in pairs by LLMs. Therefore, we also ignore the code translation task for experiments with LLMs.
II-F Implementations
All experiments for PLMs and LLMs are conducted on two Nvidia A40 GPUs with 48GB of memory for each. For pre-training of the PLMs, we perform 20 epochs with a fixed learning rate of and a batch size of 32. For fine-tuning on downstream tasks, we train for 50 epochs on both the code translation and the generation tasks, ensuring full convergence. Our fine-tuning framework is adapted from the replication package of CodeXGLUE’s [10]. When inference, we fetch the first output for evaluation.
To enable direct inference by LLM, we deploy LLaMA and StarCoder locally with the model initialization and weight-loading from Hugging Face [41]. Subsequently, we perform inference on code tasks using a series of basic prompts.
To alleviate the chaos of outputs, we additionally provide a one-shot example for in-context learning and retrieve only the first output for evaluation. The inference process is conducted with the following decoding parameters: temperature = 0.8, top_p = 0.95, and top_k = 50.
@@ II-F Implementations
To mitigate the influence of randomness and enhance the reliability of the results, all experiments concerning PLMs and LLMs are repeated five times, respectively. Besides, we employ the non-parametric Mann–Whitney U test [42] with a two-sided hypothesis to assess whether the performance differences between experimental and control groups are statistically significant. For each test, the null hypothesis is defined as: “There is no statistically significant difference between the results of the two groups.” All statistical tests are conducted at a significance level of = 0.05, corresponding to a 95% confidence level for determining statistical significance.
III Experimental Result
@@
This section discusses the experimental results targeting each RQ proposed in Section II-A. Owing to the page limit, we only showcase results with their average score and standard deviation.
III-A Results of the Input-Only Contamination Setting (RQ1)
@@
As illustrated in Section II-A1, we conduct experiments on code translation and generation tasks for PLMs while carrying out code translation and summarization tasks for LLMs. Below, we introduce their experimental results and corresponding analysis, respectively.
@@
As for PLMs, the performance of RoBERTa and GPT-2 under control and experimental groups is reported in Table II. Apparently, the performance in the contaminated experimental group does not exhibit a significant performance gain compared to the uncontaminated counterparts, regardless of the code translation or generation task and RoBERTa or GPT-2. For example, on the code translation task with GPT-2, the experimental group’s average BLEU score is only 0.97% higher than that of the control group, which falls within the margin of experimental error. Even surprisingly, the model performance of the control group surpasses that of the experimental group when experimenting with RoBERTa, although with only 0.046%–0.86% across both metrics and code tasks. Statistical tests further support this observation, with -values exceeding 0.05 among five consecutive trials, indicating no significant differences. These results suggest that the performance of PLMs will not be overestimated in the downstream tasks, even though their pre-training corpora are contaminated by the input portions of testing sets. This conclusion aligns with our expectation and is reasonable for explanation, as PLMs do not learn any mapping from input to output portions of the testing sets. Their pre-training corpora only contain questions but no answers.
@@
As for LLMs, their experimental results of the input-only contamination setting are also listed in Table II. As can be seen, the inference performance in the contaminated experimental group and the uncontaminated control group has no obvious discrepancy among the five consecutive trials. Besides, statistical tests further verify the nonsignificance of this difference with p-values all larger than 0.05. The reason is the same as that we mentioned in the PLM part, as LLMs only saw the input portions during their pre-training, considering the substantial discrepancies between inputs and outputs in code tasks, by no means can they carry any superiority in generating the output portions.
Model | Task | BLEU (Experimental Group / Control Group / p-value) | METEOR (Experimental Group / Control Group / p-value) |
RoBERTa | JavaC# | / / 0.296 | / / 0.531 |
NLJava | / / 0.210 | / / 0.095 | |
GPT-2 | JavaC# | / / 0.296 | / / 0.249 |
NLJava | / / 0.210 | / / 0.296 | |
LLaMA | JavaC# | / / 0.676 | / / 0.296 |
JavaNL | / / 0.531 | / / 0.210 | |
Starcoder | JavaC# | / / 0.296 | / / 0.403 |
JavaNL | / / 0.531 | / / 1.000 |
Note: In the Task column, the contaminated part is highlighted in bold. The results are presented in the form of , where is the average result of 5 experiments, and is the corresponding standard deviation. p-value specifies the significance of statistical tests. We highlight the higher scores between the two groups with a yellow background. Subsequent tables follow the same convention.
III-B Results of the Output-only Contamination Setting (RQ2)
@@
As elaborated in Section II-A2, we carry out experiments on code translation and generation tasks for both PLMs and LLMs. Below, we introduce their experimental results and corresponding analysis, respectively.
@@
As for PLMs, the results for RoBERTa and GPT-2 under this setting are shown in Table III. Similar to the input-only contamination experiments, there is no significant difference in performance between the contaminated experimental group and the uncontaminated control group, regardless of any code intelligence task and PLMs under test. Specifically, experiments with RoBERTa all demonstrate a slightly superior performance in the contaminated scenarios, while GPT-2 performs the opposite. However, the differences are minimal. For example, in terms of BLEU score on the code translation task, the experimental group of Roberta was only 0.21% higher than the control group, and the control group of GPT-2 was just 0.51% higher than the experimental group on average. Besides, when focusing on statistical tests, their results among the five trials do not manifest any statistical significance with p-values all larger than 0.05. This indicates that exposure to the output portions of the testing set during the pre-training stage does not cause performance overestimation for PLMs.
@@
As for LLMs, the inference results under the output-only contamination setting are presented in Table III as well. The observed performance distribution aligns with that of PLMs, i.e., neither the contaminated experimental group nor the uncontaminated control group dominate on code intelligence tasks, and none of the statistical tests demonstrate significant differences with p-values larger than 0.05 consistently. This suggests that contamination appearing on output portions of testing sets will not induce performance overestimation in downstream tasks.
In summary, neither PLMs nor LLMs will be influenced by the output-only contamination in their pre-training corpora. As we explained in Section III-A, models only saw the partial samples of testing sets, which is not helpful for them to infer the rest of the samples, as the differences between input and output portions of code intelligence tasks are substantial.
Model | Task | BLEU (Experimental Group / Control Group / p-value) | METEOR (Experimental Group / Control Group / p-value) |
RoBERTa | JavaC# | / / 0.404 | / / 0.095 |
NLJava | / / 0.144 | / / 0.676 | |
GPT-2 | JavaC# | / / 0.531 | / / 0.676 |
NLJava | / / 0.600 | / / 0.403 | |
LLaMA | C#Java | / / 0.676 | / / 1.000 |
NLJava | / / 0.403 | / / 0.210 | |
Starcoder | C#Java | / / 0.144 | / / 0.403 |
NLJava | / / 0.210 | / / 0.095 |
Model | Task | BLEU (Experimental Group / Control Group / p-value) | METEOR (Experimental Group / Control Group / p-value) |
RoBERTa | JavaC# | / / 0.210 | / / 0.296 |
NLJava | / / 0.531 | / / 0.144 | |
GPT-2 | JavaC# | / / 1.000 | / / 0.676 |
NLJava | / / 0.835 | / / 0.531 | |
LLaMA | JavaC# | / / 0.144 | / / 0.095 |
C#Java | / / | / / 0.060 | |
Starcoder | JavaC# | / / 0.676 | / / 1.000 |
C#Java | / / | / / 0.095 |
Note: Hypothesis test values less than 0.05 are indicated in red. Subsequent tables follow the same convention.
III-C Results of the Unpaired Contamination Setting (RQ3)

### v3

@@ 2.5.3. RQ3: Does Unpaired Contamination Lead to Performance Overestimation?
(2) LLMs (Pre-train, then infer): Constructing contaminated experimental groups in this RQ is relatively difficult compared with previous ones, as it requires both input and output portions of testing sets explicitly existing in the pre-training corpora of LLMs, but as independent samples. For code generation and summarization tasks, although finding independent code snippets in LLMs’ pre-training corpora is easy, it is challenging to find an independent NL that coincidentally describes a code functionality but is not attached to a code snippet, because code and comments are normally closely attached in the wild. Therefore, we ignore the above two tasks for LLMs in the unpaired contamination setting and focus on the code translation task only.
Considering that manually finding those PL pairs (e.g., Java-C# and Python-Java) in the tremendous pre-training corpora is extremely time-consuming, we propose a series of filtering rules to substantially shrink the searching range. 1) Functions of different PLs with identical function names normally carry the same functionalities. Thus, we first use tree-sitter to parse their function names, then remove special symbols (e.g., _) to keep letters only, and switch them all to lowercase for the initial cross-PL function pair construction. 2) For each pair of cross-PL functions, we count some key features from their Abstract Syntax Trees (ASTs) for complexity comparison, thereby proceeding with further filtering. To be specific, we count their total number of AST nodes, statement nodes, expression nodes, and function-call nodes as key features. Function pairs on the ratios of all kinds of node counts within the range of are kept; otherwise, they are discarded. 3) For the remaining function pairs, we extract their keywords (e.g., while, if, and +) and compute the Generalized Jaccard Index (GJI) (Zhang et al., 2019a) as below for semantic-level filtering:
where and denote the source and target functions, represents the monitored keyword set, and is the count of keyword in the function or . Function pairs with are kept. 4) Finally, we further conduct an abstractive execution flow comparison based on ASTs. In detail, we extract the AST nodes in 2) with pre-order traversal to construct node sequences and compute the Longest Common Subsequence (LCS) (Bergroth et al., 2000) between function pairs:
where and are the node sequences of the source and target functions, and denotes the sequence length. Function pairs with are kept.
Initially, we obtain over 60M Python-Java and 90M Java-C# function pairs in Google BigQuery and 29M and 40M corresponding function pairs in Stack after the above step 1). After the following automated filtering steps 2)-4), we retain a much smaller and higher-quality candidate pool with 50,952 Python-Java and 60,900 Java-C# function pairs from Stack, while 29,318 and 13,462 corresponding function pairs from Google BigQuery for follow-up manual screening. As can be seen, the manual effort is substantially reduced. Ultimately, for each code translation task, we manually select 100 function pairs to construct the contaminated experimental group. As for the uncontaminated control group construction and subsequent experiments, we follow the same procedure as previous RQs. Detailed statistics are shown in Table 1.
2.5.4. RQ4: Does Paired Contamination Lead to Performance Overestimation?
(1) PLMs (Pre-train, fine-tune, then infer): Regarding the Java side, code translation task (JavaC#) under this setting continues following all the procedures described in RQ1. The key difference lies in the construction of the contaminated experimental group. We concatenate the input (i.e., Java) and output (i.e., C#) portions of each sample in the CodeTrans test set using a separator token <SEP>, and inject these concatenated samples into the pre-training corpora of PLMs. We adopt <SEP> to distinguish different PLs, due to its wide usage in practice for bimodal pre-training (Feng et al., 2020; Wang et al., 2021; Zhu et al., 2024). The same procedures are applied to the code generation and summarization tasks of the Java side as well as to the corresponding Python-related experiments.
(2) LLMs (Pre-train, then infer): On the Java side, to construct the contaminated experimental group for the code generation task (NLJava) under this setting, we examine each Java sample in BigQuery and Stack using tree-sitter to determine whether the NL description appears immediately before the function signature. After that, we perform manual filtering to select 100 samples whose NL descriptions are accurate and informative enough to be used for code generation. For the code summarization task (JavaNL) under this setting, a paired contaminated sample should be an ordered concatenation of NL and Java code. However, it is rare to find a comment that appears after the function body both in the wild and LLMs’ pre-training corpora. Therefore, we instead extract comments following the function signatures as desired NLs via tree-sitter, along with these functions, to construct the paired contamination samples. The reasons are two-fold: 1) Such comments often encapsulate the summarization of the function. 2) Such comments lie in front of their corresponding function body, largely satisfying our predefined settings. We then manually inspect the samples to determine whether the comments can serve as valid summaries, curating 100 proper samples for the experimental group. The uncontaminated control group construction of the above two code tasks and their follow-up experiments are carried out following the same procedures described in RQ1. The aforementioned process is consistently applied to the Python-related tasks.
In practice, code snippets with equivalent functionalities but written in different PLs are rarely placed in the same source files. As a result, it is difficult to find cross-PL program pairs with equivalent logic that appear as a sample in the pre-training corpora of LLMs. For this reason, we ignore the code translation task for both the Java and Python sides in the experiments with LLMs.
2.6. Implementations
For pre-training of the PLMs, we perform 20 epochs with a fixed learning rate of and a batch size of 32. For fine-tuning on downstream tasks, we train each model for 50 epochs on all three tasks to ensure full convergence, and select the model that achieves the best performance on the validation set. Our fine-tuning framework is adapted from the replication package of CodeXGLUE’s (Lu et al., 2021). When inference, we fetch the first output for evaluation. To enable direct inference by LLM, we deploy LLaMA and StarCoder locally with the model initialization and weight loading from Hugging Face (Wolf et al., 2020). Subsequently, we perform inference on code tasks using a series of basic prompts.To alleviate the chaos of outputs, we additionally provide a one-shot example for in-context learning and retrieve only the first output for evaluation. The inference process is conducted with the following decoding parameters: temperature = 0.1, top_p = 0.95, and top_k = 50.
@@ 2.6. Implementations
To mitigate the influence of randomness and enhance the reliability of the results, all experiments concerning PLMs and LLMs are repeated five times, respectively. Besides, we employ the non-parametric Mann–Whitney U test (Mann and Whitney, 1947) with a two-sided hypothesis to assess whether the performance differences between experimental and control groups are statistically significant. For each test, the null hypothesis is defined as: “There is no statistically significant difference between the results of the two groups.” All statistical tests are conducted at a significance level of = 0.05, corresponding to a 95% confidence level for determining statistical significance. All experiments above are conducted on six Nvidia A40 GPUs with 48GB of memory for each. The pre-training, fine-tuning, and inference of two PLMs across three code tasks and two PLs require 3,500 GPU-hours, while the inference of two LLMs for all experiments requires 100 GPU-hours.
3. Experimental Results
@@
This section discusses the experimental results targeting each RQ proposed in Section 2.1.
| Metric | Model Performance (Experimental Group / Control Group / p-value) | |||
| RoBERTa | GPT-2 | LLaMA | StarCoder | |
| Code Translation: Java C# | ||||
| BLEU | / 86.11±0.08 / 0.296 | 37.52±0.58 / / 0.296 | / 72.11±1.16 / 0.676 | / 73.70±1.62 / 0.296 |
| METEOR | / 88.99±0.15 / 0.531 | 48.13±0.51 / / 0.249 | / 87.72±0.79 / 0.296 | / 85.41±1.54 / 0.403 |
| Code Translation: Python Java | ||||
| BLEU | 70.08±0.30 / 69.78±0.22 / 0.296 | 48.73±0.09 / 48.67±0.13 / 0.526 | 28.55±0.17 / 28.22±0.22 / 0.296 | / 37.69±0.65 / 0.530 |
| METEOR | 70.32±0.34 / 69.97±0.22 / 0.095 | 49.70±0.09 / 49.57±0.12 / 0.144 | 44.91±0.16 / 44.86±0.27 / 0.835 | 50.63±0.40 / 50.29±0.75 / 0.531 |
| Code Generation: NL Java | ||||
| BLEU | / 40.77±0.42 / 0.210 | 17.99±0.31 / / 0.210 | —— / —— / —— | —— / —— / —— |
| METEOR | / 51.47±0.38 / 0.095 | 22.26±0.33 / / 0.296 | —— / —— / —— | —— / —— / —— |
| Code Generation: NL Python | ||||
| BLEU | 37.46±0.23 / 37.39±0.13 / 0.834 | / 15.47±0.09 / 0.173 | —— / —— / —— | —— / —— / —— |
| METEOR | 45.36±0.28 / 45.31±0.09 / 0.835 | / 31.63±0.15 / 0.144 | —— / —— / —— | —— / —— / —— |
| Code Summarization: Java NL | ||||
| BLEU | 39.24±0.28 / 38.68±0.42 / 0.296 | / 2.80±0.06 / 0.249 | / 10.95±0.92 / 0.531 | 9.99±0.48 / / 0.531 |
| METEOR | 48.85±0.41 / 48.51±0.34 / 0.249 | / 18.72±0.22 / 0.600 | / 30.79±1.02 / 0.210 | 28.51±0.44 / / 1.000 |
| Code Summarization: Python NL | ||||
| BLEU | 37.39±0.21 / 37.09±0.25 / 0.144 | / 7.15±0.03 / 0.401 | 10.17±0.43 / 9.82±0.26 / 0.296 | 6.98±0.33 / 6.75±0.21 / 0.296 |
| METEOR | 46.60±0.32 / 46.32±0.16 / 0.172 | / 22.12±0.10 / 0.295 | / 32.71±0.63 / 0.210 | / 27.33±0.17 / 0.144 |
Note: Owing to the page limit, each cell only showcases results with their average score and standard deviation of experimental and control groups, respectively, as well as corresponding statistical tests (p-value). In addition, p-values less than 0.05 are indicated in red. The subsequent tables follow the same convention.
3.1. Results of the Input-Only Contamination Setting (RQ1)
@@
As illustrated in Section 2.1.1, we conduct experiments on code translation, generation, and summarization tasks for PLMs while carrying out code translation and summarization tasks for LLMs. Below, we introduce their experimental results and corresponding analysis, respectively.
@@
As for PLMs (Pre-train, fine-tune, then infer), the performance of RoBERTa and GPT-2 under control and experimental groups is reported in Table 2. Apparently, the performance in the contaminated experimental group does not exhibit a significant performance gain compared to the uncontaminated counterparts, regardless of any code task, PLMs, and PLs, always fluctuating in the range of (-0.85%)–1.45%. For example, on the code translation task with GPT-2, the experimental group’s average BLEU scores on both PLs are only 0.12%–0.97% higher than those of the control group. As for RoBERTa, the stable gain of contamination is limited to the code summarization task, but only surpassing 0.6%–1.45% to the uncontaminated control group, showing that the above all falls within the margin of experimental error. Even surprisingly, the model performance of the experimental group underperforms that of the control group when experimenting with GPT-2 on the code summarization task, although with only a lagging of (-1.07%)–(-0.21%) across both metrics and PLs. Statistical tests across all three code tasks further support this observation, with -values exceeding 0.05 among five consecutive trials, indicating no significant differences. These results suggest that the performance of PLMs will not be overestimated in the downstream tasks, even though their pre-training corpora are contaminated by the input portions of testing sets. This conclusion aligns with our expectation and is reasonable for explanation, as PLMs do not learn any mapping from input to output portions of the testing sets. Their pre-training corpora only contain questions, but no answers.
@@
As for LLMs (Pre-train, then infer), their experimental results of the input-only contamination setting are also listed in Table 2. As can be seen, the inference performance of LLaMA and StarCoder in the contaminated experimental group and the uncontaminated control group still has no obvious discrepancy among the five consecutive trials, manifesting a relative performance difference of only (-3.12%)–3.56% across both code translation and summarization tasks. Besides, statistical tests further verify the nonsignificance of this difference with p-values all larger than 0.05. The reason is the same as that we mentioned in the PLM part, as LLMs only saw the input portions during their pre-training, considering the substantial discrepancies between inputs and outputs in code tasks, by no means can they carry any superiority in generating the output portions.
3.2. Results of the Output-only Contamination Setting (RQ2)
@@
As elaborated in Section 2.1.2, PLMs are still experimented on all studied code tasks, while LLMs are assessed on two of the code tasks, i.e., code translation and generation. Below, we introduce their experimental results and corresponding analysis, respectively.
@@
As for PLMs (Pre-train, fine-tune, then infer), the results for RoBERTa and GPT-2 under this setting are shown in Table 3. Similar to the input-only contamination experiments, there is no significant difference in performance between the contaminated experimental group and the uncontaminated control group, regardless of any code task and PLMs under test. Specifically, experiments with RoBERTa demonstrate a slightly superior performance of 0.07%–0.93% and 0.02%–0.87% in terms of BLEU and METEOR under the contaminated scenarios of almost all code tasks, while GPT-2 performs the opposite, manifesting a consistent but slight decline of (-1.07%)–(-0.02%) and (-1.45%)–(-0.02%) in terms of each metric under the contaminated scenario of all code tasks. However, the difference is minimal, and when focusing on statistical tests, the results among their corresponding five trials do not manifest any statistical significance with p-values all larger than 0.05. This indicates that exposure to the output portions of the testing set during the pre-training stage does not cause significant performance overestimation for PLMs.
@@
As for LLMs (Pre-train, then infer), the inference results under the output-only contamination setting are presented in Table 3 as well. As can be seen, the contaminated experimental groups of LLaMA and StarCoder indeed perform slightly better overall, but the superiority is restricted to 0.3%–1.09% and (-0.47%)–1.46% in terms of BLEU and METEOR across both code tasks for the former, as well as 0.05%–2.89% and (-0.91%)–1.46% for the latter. In addition, none of the statistical tests demonstrate significant differences with p-values larger than 0.05 consistently. This suggests that contamination appearing on output portions of testing sets will not induce significant performance overestimation in downstream tasks. As we explained in Section 3.1, models only saw the partial samples of testing sets, which is not helpful for them to infer the rest of the samples, as the differences between input and output portions of code tasks are substantial.
| Metric | Model Performance (Experimental Group / Control Group / p-value) | |||
| RoBERTa | GPT-2 | LLaMA | StarCoder | |
| Code Translation: Java C# | ||||
| BLEU | 86.29±0.45 / 86.11±0.09 / 0.404 | 36.97±0.31 / 37.16±0.36 / 0.531 | 54.09±0.11 / 53.93±0.15 / 0.116 | 59.64±0.19 / 59.61±0.06 / 0.676 |
| METEOR | 89.17±0.13 / 88.99±0.16 / 0.095 | 47.66±0.35 / 47.75±0.32 / 0.676 | 74.76±0.10 / 74.40±0.23 / 0.075 | 79.89±0.46 / 79.47±0.32 / 0.144 |
| Code Translation: Python Java | ||||
| BLEU | 69.83±0.08 / 69.78±0.22 / 0.917 | 48.66±0.14 / 48.67±0.13 / 0.834 | 48.88±0.72 / 48.81±0.39 / 0.835 | 53.83±0.30 / 53.46±0.62 / 0.296 |
| METEOR | 70.21±0.14 / 69.97±0.22 / 0.142 | 49.56±0.11 / 49.57±0.12 / 0.916 | 62.94±0.78 / 62.09±0.33 / 0.095 | 74.13±0.25 / 74.81±0.35 / 0.095 |
| Code Generation: NL Java | ||||
| BLEU | 41.07±1.08 / 40.77±0.43 / 0.144 | 17.63±0.14 / 17.72±0.20 / 0.600 | 48.20±1.07 / 47.68±0.87 / 0.403 | 47.65±1.11 / 46.31±0.96 / 0.210 |
| METEOR | 51.49±0.87 / 51.47±0.38 / 0.676 | 21.95±0.264 / 21.98±0.13 / 0.403 | 62.37±0.91 / 61.47±0.93 / 0.210 | 64.45±0.88 / 63.52±0.70 / 0.095 |
| Code Generation: NL Python | ||||
| BLEU | 37.60±0.25 / 37.39±0.13 / 0.209 | 15.36±0.22 / 15.47±0.09 / 0.600 | 30.58±0.77 / 30.32±0.50 / 0.531 | 49.19±0.38 / 49.01±0.23 / 0.531 |
| METEOR | 45.46±0.23 / 45.31±0.09 / 0.531 | 31.17±0.10 / 31.63±0.15 / 0.116 | 44.77±0.90 / 44.98±0.62 / 1.000 | 62.00±0.37 / 61.82±0.63 / 0.403 |
| Code Summarization: Java NL | ||||
| BLEU | 39.04±0.66 / 38.68±0.42 / 1.000 | / 2.80±0.06 / 0.528 | —— / —— / —— | —— / —— / —— |
| METEOR | 48.93±0.57 / 48.51±0.34 / 0.249 | / 18.72±0.22 / 0.917 | —— / —— / —— | —— / —— / —— |
| Code Summarization: Python NL | ||||
| BLEU | / 37.09±0.25 / 0.676 | / 7.15±0.03 / 0.399 | —— / —— / —— | —— / —— / —— |
| METEOR | 46.33±0.13 / 46.32±0.16 / 1.000 | / 22.12±0.10 / 0.753 | —— / —— / —— | —— / —— / —— |
| Metric | Model Performance (Experimental Group / Control Group / p-value) | |||
| RoBERTa | GPT-2 | LLaMA | StarCoder | |
| Code Translation: Java C# | ||||
| BLEU | 86.22±0.22 / 86.11±0.09 / 0.210 | 37.15±0.37 / 37.16±0.36 / 1.000 | 66.93±1.15 / 65.73±1.29 / 0.144 | 76.72±1.25 / 76.17±1.45 / 0.676 |
| METEOR | 89.09±0.26 / 88.99±0.16 / 0.296 | 47.82±0.18 / 47.75±0.32 / 0.676 | 80.88±0.97 / 79.56±0.55 / 0.095 | 88.85±1.11 / 88.89±1.07 / 1.000 |
| Code Translation: Python Java | ||||
| BLEU | 70.14±0.32 / 69.78±0.22 / 0.060 | 48.70±0.06 / 48.67±0.13 / 0.834 | 28.11±0.14 / 27.75±0.25 / 0.059 | 33.06±0.43 / 32.91±0.30 / 0.530 |
| METEOR | 70.42±0.39 / 69.97±0.22 / 0.095 | 49.61±0.09 / 49.57±0.12 / 0.835 | 39.21±0.07 / 39.36±0.23 / 0.346 | 45.89±0.57 / 46.73±0.37 / 0.095 |
| Code Generation: NL Java | ||||
| BLEU | 40.98±0.54 / 40.77±0.43 / 0.531 | 17.73±0.25 / 17.72±0.20 / 0.835 | —— / —— / —— | —— / —— / —— |
| METEOR | 51.74±0.18 / 51.47±0.38 / 0.144 | 22.13±0.35 / 21.98±0.13 / 0.531 | —— / —— / —— | —— / —— / —— |
| Code Generation: NL Python | ||||
| BLEU | 37.59±0.19 / 37.39±0.13 / 0.059 | 15.36±0.08 / 15.47±0.09 / 0.095 | —— / —— / —— | —— / —— / —— |
| METEOR | 45.46±0.21 / 45.31±0.09 / 0.296 | 31.36±0.27 / 31.63±0.15 / 0.095 | —— / —— / —— | —— / —— / —— |
| Code Summarization: Java NL | ||||
| BLEU | / 38.68±0.42 / 0.296 | 2.81±0.03 / 2.80±0.06 / 1.000 | —— / —— / —— | —— / —— / —— |
| METEOR | / 48.51±0.34 / 0.917 | / 18.72±0.22 / 0.531 | —— / —— / —— | —— / —— / —— |
| Code Summarization: Python NL | ||||
| BLEU | / 37.09±0.25 / 0.835 | / 7.15±0.03 / 0.142 | —— / —— / —— | —— / —— / —— |
| METEOR | / 46.32±0.16 / 0.462 | / 22.12±0.10 / 0.531 | —— / —— / —— | —— / —— / —— |
3.3. Results of the Unpaired Contamination Setting (RQ3)